# Utilizing Large Language Models in an iterative paradigm with domain feedback for zero-shot molecule optimization

### v7

@@ Abstract
Molecule optimization is a critical task in drug discovery to optimize desired properties of a given molecule through chemical modification. Despite Large Language Models (LLMs) holding the potential to efficiently simulate this task by using natural language to direct the optimization, straightforwardly utilizing them shows limited performance. In this work, we facilitate utilizing LLMs in an iterative paradigm by proposing a simple yet highly effective domain feedback provider, namely $\text{Re}^{3}$. In detail, $\text{Re}^{3}$ harnesses an external toolkit, RDKit, to handle the molecule hallucination, if the modified molecule is chemically invalid. Otherwise, its desired properties are computed and compared to the original one, establishing reliable domain feedback with correct direction and distance towards the objective, followed by a retrieved example, to guide the LLM to refine the modified molecule. We conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{3}$ shows significant improvements. Particularly, for 20 single-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 16.96% and 20.76% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 6.04% and 5.25%.
@@
Khiem Le, Nitesh V. Chawla Department of Computer Science and Engineering University of Notre Dame, IN, USA {kle3, nchawla}@nd.edu
@@ 1 Introduction
Drug discovery, aimed at identifying molecules that can safely modulate the disease Bateman (2022), is a costly process that takes on average more than a decade to reach approval and release Berdigaliyev and Aljofan (2020). Fortunately, recent advancements in AI for molecular learning have been revolutionizing this process by remarkably accelerating many of its molecule-related tasks. While tasks such as molecule property prediction, as well as synthesis and retrosynthesis have been extensively investigated Zhao et al. (2023); Shi et al. (2023); Liu et al. (2024b); Le et al. (2024), molecule optimization, a critical task, remains largely under-explored.
Figure 1: An illustration of Refinement with $\text{Re}^{3}$ and how the procedure differs from the previous ReDF.
@@
Molecule optimization is the task of modifying a given molecule with the objective of achieving desired properties. Importantly, the modified molecule needs to retain a certain level of similarity to the original one, known as the similarity constraint of the modified molecule Jorgensen (2009). This task is routine in pharmaceutical companies and conventionally carried out by chemists through manual attempts Hoffer et al. (2018); de Souza Neto et al. (2020), which is not scalable. Given this context, the latest groundbreaking in AI, Large Language Models (LLMs), hold tremendous potential to efficiently simulate and scale this task by using natural language to direct the optimization Zhang et al. (2024). Unlike the traditional supervised training approach Loeffler et al. (2024); Xia et al. (2024), utilizing LLMs with outstanding abilities alleviates the need for labeled molecules Liu et al. (2023a), as well as exhibits zero-shot and open-vocabulary generalization beyond a pre-defined set of objectives, making them highly promising for this task.
@@
As pioneers, Zhang et al. (2024) straightforwardly explore the capability of LLMs for solving this task using natural language to direct the optimization and notice limited performance. Moving further, motivated by the iterative refinement strategy employed by chemists, Liu et al. (2024a) propose to utilize LLMs in an iterative paradigm with domain feedback. Specifically, a domain feedback provider, ReDF verifies whether the modified molecule meets the objective, if not, ReDF retrieves from a prior database a molecule that is similar to the modified molecule and meets the objective, serving as an example to guide the LLM to refine the modified molecule. Notably, the domain feedback provider plays a vital role in the iterative paradigm Yang et al. (2022); Madaan et al. (2024). Although bringing improved performance, ReDF does not consider the molecule hallucination where the modified molecule is chemically invalid Guo et al. (2023), preventing the retrieving step. Moreover, by solely depending on a retrieved example as domain feedback, ReDF unexpectedly harms the similarity constraint of the modified molecule.
@@
In this work, we propose an improved domain feedback provider, namely $\text{Re}^{3}$, which is carefully designed to facilitate utilizing LLMs in an iterative paradigm more effectively, as illustrated in Figure 1. Different from the previous ReDF, first of all, $\text{Re}^{3}$ takes into account the molecule hallucination, thereby actively promoting the retrieving step. To be more precise, at each iteration, the modified molecule is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013). If the modified molecule is chemically invalid, the concrete error message from RDKit such as the existence of an unclosed ring Schoenmaker et al. (2023); Skinnider (2024), is extracted as reliable domain feedback to inform the LLM to correct the modified molecule, in analogy to debugging Chen et al. (2023); Zhong et al. (2024). Otherwise, $\text{Re}^{3}$ verifies whether the modified molecule meets the objective, if not, its desired properties are computed and compared to the original one, establishing reliable domain feedback with correct direction and distance towards the objective to explicitly guide the LLM to refine the modified molecule. Additionally, as in the previous ReDF, a retrieved example is appended to further enrich the domain feedback. By diminishing dependence on a retrieved example, $\text{Re}^{3}$ excels at satisfying the similarity constraint of the modified molecule, as we demonstrate along with improved performance.
@@
We first revisit the evaluation metric used in previous works Liu et al. (2023b, 2024a), Hit ratio as the percentage of successful modifications, and identify a significant flaw in that failures involving modified molecules that are chemically invalid are excluded in calculating Hit ratio, resulting in an artificially high value. After rectifying, we conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{3}$ shows significant improvements over baselines and excels at satisfying the similarity constraint of the modified molecule. Particularly, for 20 single-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 16.96% and 20.76% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{3}$ enhances Hit ratio by 6.04% and 5.25%.
@@ 2 Methodology
Here we delve into the details of utilizing LLMs in an iterative paradigm integrated with the proposed domain feedback provider, $\text{Re}^{3}$.
@@
Algorithm 1 Refinement with $\text{Re}^{3}$
@@ 2.1 Preliminaries
where ||||| | indicates concatenation, $E(C_{p}(\widehat{m}))\in\{0,1\}$ indicates whether $\widehat{m}$ meets the objective with its desired properties $C_{p}(\widehat{m})$.
@@
vanilla Self-feedback ReDF (3) Re³DF (3) vanilla Self-feedback ReDF (3) Re³DF (3)
@@
2.2 Refinement with $\text{Re}^{3}$
@@
After initialization, $\widehat{m}$ is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013), a widely-used cheminformatics toolkit. If $\widehat{m}$ is chemically invalid, the concrete error message from RDKit is extracted as reliable domain feedback to inform the LLM $\mathcal{M}$ to correct $\widehat{m}$. Otherwise, $\text{Re}^{3}$ verifies whether $\widehat{m}$ meets the objective by checking whether $E(C_{p}(\widehat{m}))=1$, if not, its desired properties $C_{p}(\widehat{m})$, are computed and compared to the original one, establishing reliable domain feedback, denoted fb, with correct direction and distance towards the objective, $|C_{p}(\widehat{m})-C_{p}(m)|$, to explicitly guide the LLM $\mathcal{M}$ to refine $\widehat{m}$. Additionally, $\text{Re}^{3}$ retrieves from a prior database $\mathcal{D}$ a molecule $m_{e}$ that is similar to $\widehat{m}$ and meets the objective, serving as an example to further enrich the domain feedback:
@@
where $\left\langle m_{e}^{\prime},\widehat{m}\right\rangle$ is the Tanimoto similarity Bajusz et al. (2015) between a molecule $m_{e}^{\prime}$ in $\mathcal{D}$ and $\widehat{m}$. Subsequently, the described procedure is iterated for a pre-defined maximum number of iterations N or either can be stopped early if $\widehat{m}$ meets the objective, as precisely depicted in Algorithm 1.

### v8

@@ Abstract
Molecule optimization is a critical task in drug discovery to optimize desired properties of a given molecule. Despite Large Language Models (LLMs) holding the potential to efficiently simulate this task by using natural language to direct the optimization, straightforwardly utilizing them shows limited performance. In this work, we facilitate utilizing LLMs in an iterative paradigm by proposing a simple yet effective domain feedback provider, namely $\text{Re}^{2}$. In detail, $\text{Re}^{2}$ harnesses an external toolkit, RDKit, to handle the molecule hallucination, if the modified molecule is chemically invalid. Otherwise, $\text{Re}^{2}$ verifies whether the modified molecule meets the objective, if not, its desired properties are computed and compared to the original one, establishing reliable domain feedback with correct direction and distance towards the objective to explicitly guide the LLM to refine the modified molecule. We conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{2}$ shows significant improvements. Notably, for 20 single-property objectives, $\text{Re}^{2}$ enhances Hit ratio by 16.96% and 20.76% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{2}$ enhances Hit ratio by 6.04% and 5.25%.
@@
Khiem Le, Nitesh V. Chawla University of Notre Dame, IN, USA {kle3, nchawla}@nd.edu Github: https://github.com/lhkhiem28/Re2DF
@@ 1 Introduction
Drug discovery, aimed at identifying molecules that can safely modulate the disease Bateman (2022), is a costly process that takes on average more than a decade to reach approval and release Berdigaliyev and Aljofan (2020). Fortunately, recent advancements in AI for molecular science have been revolutionizing this process by remarkably accelerating many of its molecule-related tasks. While tasks such as molecule property prediction, as well as synthesis and retrosynthesis have been extensively investigated Zhao et al. (2023); Shi et al. (2023); Liu et al. (2024b); Le et al. (2024), molecule optimization, a critical task, remains largely under-explored.
Figure 1: An illustration of Refinement with $\text{Re}^{2}$. The domain feedback is generated from $\text{Re}^{2}$ to guide the LLM to refine the modified molecule.
@@
Molecule optimization is the task of modifying a given molecule with the objective of achieving desired properties. Importantly, the modified molecule needs to retain a certain level of similarity to the original one, known as the similarity constraint of the modified molecule Jorgensen (2009). This task is routine in pharmaceutical companies and conventionally carried out by chemists through manual attempts Hoffer et al. (2018); de Souza Neto et al. (2020), which is not scalable. Given this context, the latest groundbreaking in AI, Large Language Models (LLMs), hold tremendous potential to efficiently simulate and scale this task by using natural language to direct the optimization Zhang et al. (2024). Unlike the supervised fine-tuning approach Loeffler et al. (2024); Xia et al. (2024), utilizing LLMs with outstanding abilities alleviates the need for labeled molecules Liu et al. (2023a), as well as exhibits zero-shot and open-vocabulary generalization beyond a pre-defined set of objectives, making them promising for this task.
@@
As pioneers, Zhang et al. (2024) straightforwardly explore the capability of LLMs for solving this task by using natural language to direct the optimization and notice limited performance. Moving further, motivated by the iterative refinement strategy employed by chemists, Liu et al. (2024a) propose to utilize LLMs in an iterative paradigm with domain feedback. Specifically, a domain feedback provider, ReDF verifies whether the modified molecule meets the objective, if not, ReDF retrieves from a prior database a molecule that is similar to the modified molecule and meets the objective, serving as an example to guide the LLM to refine the modified molecule. Although bringing improved performance, ReDF does not consider the molecule hallucination where the modified molecule is chemically invalid Guo et al. (2023), preventing the retrieving step. Moreover, by solely depending on a retrieved example as domain feedback, ReDF unexpectedly harms the similarity constraint of the modified molecule.
@@
In this work, we propose an improved domain feedback provider, namely $\text{Re}^{2}$, which is carefully designed to facilitate utilizing LLMs in an iterative paradigm more effectively, illustrated in Figure 1. Different from the previous ReDF, $\text{Re}^{2}$ takes into account the molecule hallucination, thereby actively promoting the retrieving step. To be more precise, at each iteration, the modified molecule is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013). If the modified molecule is chemically invalid, the concrete error message from RDKit such as the existence of an unclosed ring Schoenmaker et al. (2023); Skinnider (2024), is extracted as reliable domain feedback to inform the LLM to correct the modified molecule, in analogy to debugging Chen et al. (2023); Zhong et al. (2024). Otherwise, $\text{Re}^{2}$ verifies whether the modified molecule meets the objective, if not, its desired properties are computed and compared to the original one, establishing reliable domain feedback with correct direction and distance towards the objective to explicitly guide the LLM to refine the modified molecule. Additionally, as the previous ReDF, $\text{Re}^{2}$ retrieves from a prior database a molecule that is similar to the modified molecule and meets the objective, serving as an example to enrich the domain feedback. By diminishing dependence on a retrieved example, $\text{Re}^{2}$ excels at satisfying the similarity constraint of the modified molecule, as we demonstrate along with improved performance.
@@
We revisit the evaluation metric used in previous works Liu et al. (2023b, 2024a), Hit ratio which is the percentage of successful modifications, and identify a significant flaw in that failures involving modified molecules that are chemically invalid are excluded in calculating Hit ratio, resulting in an artificially high value. After rectifying, we conduct experiments across both single- and multi-property objectives with 2 thresholds, where $\text{Re}^{2}$ shows significant improvements over baselines and excels at satisfying the similarity constraint of the modified molecule. Notably, for 20 single-property objectives, $\text{Re}^{2}$ enhances Hit ratio by 16.96% and 20.76% under loose (l) and strict (s) thresholds, respectively. For 32 multi-property objectives, $\text{Re}^{2}$ enhances Hit ratio by 6.04% and 5.25%.
@@ 2 Methodology
Here we delve into the details of utilizing LLMs in an iterative paradigm integrated with the proposed domain feedback provider, $\text{Re}^{2}$.
@@
Algorithm 1 Refinement with $\text{Re}^{2}$
@@ 2.1 Preliminaries
where ||||| | signifies concatenation, $E(C_{p}(\widehat{m}))\in\{0,1\}$ indicates whether $\widehat{m}$ meets the objective with its desired properties $C_{p}(\widehat{m})$.
Figure 2:
@@
vanilla Self-DF (3) ReDF (3) Re²DF (3) vanilla Self-DF (3) ReDF (3) Re²DF (3)
@@
Desired properties thresholds Llama-3.1-8B Llama-3.1-70B
ReDF (3) Re²DF (3) w/o $\text{fb}_{\text{validity}}$ w/o $m_{e}$ ReDF (3) Re²DF (3) w/o $\text{fb}_{\text{validity}}$ w/o $m_{e}$
+ LogP l 58.00 61.50 80.50 82.50 64.00 65.00 74.50 78.50 78.50 78.50 92.50 93.50 75.00 75.50 93.00 94.00
s 56.50 65.50 76.50 80.00 53.50 54.50 62.50 73.00 73.00 74.00 91.50 94.00 80.50 81.00 84.00 87.50
- LogP l 56.00 72.00 77.50 84.50 63.50 68.50 46.50 75.00 85.50 85.50 98.00 98.50 83.00 83.00 85.00 90.50
s 37.00 69.50 77.00 86.50 49.00 56.50 26.50 71.00 75.50 76.50 91.00 94.50 81.00 81.50 78.00 87.50
+ TPSA l 54.00 61.00 76.00 80.00 55.50 55.50 63.50 73.50 79.00 79.50 88.50 88.50 79.00 79.00 91.00 91.00
s 44.00 54.50 72.00 77.50 57.00 59.00 59.00 76.00 76.00 76.00 92.50 92.50 73.00 73.00 88.50 88.50
- TPSA l 53.50 66.50 85.50 88.00 65.50 66.50 43.50 75.50 71.00 72.00 92.50 96.00 80.50 81.00 63.50 86.00
s 43.50 66.50 84.50 89.50 68.50 73.50 38.50 78.50 72.50 74.50 93.50 95.50 81.00 81.00 56.50 83.50
+ HBDs l 49.50 60.00 75.50 81.50 56.50 59.50 57.00 74.50 78.50 78.50 93.00 93.50 77.50 77.50 90.00 91.00
s 35.50 52.50 69.00 77.50 47.50 49.50 30.00 66.50 62.00 63.50 78.50 82.00 69.50 71.00 72.50 80.00
- HBDs l 48.98 73.47 79.59 87.76 69.57 69.57 33.33 73.91 46.94 57.14 83.67 87.76 66.67 66.67 53.62 72.46
s 28.57 57.14 75.51 83.67 56.52 69.57 10.14 71.01 38.78 51.02 79.59 83.67 66.67 69.57 11.59 75.36
+ HBAs l 52.50 63.00 73.50 81.50 58.00 60.50 58.50 72.00 75.50 75.50 87.50 88.50 78.00 78.00 88.50 89.00
s 40.00 53.50 69.00 77.00 46.00 47.50 47.00 67.50 59.50 60.50 81.00 84.00 71.50 72.00 74.00 75.50
- HBAs l 63.27 63.27 93.88 97.96 73.91 78.26 49.28 79.71 75.51 79.59 89.80 93.88 65.22 66.67 50.72 81.16
s 46.94 69.39 81.63 89.80 65.22 66.67 20.29 76.81 71.43 75.51 83.67 89.80 65.22 71.01 20.29 78.26
+ Druglikeness l 46.50 67.50 74.50 84.00 54.50 61.50 39.00 79.00 64.50 78.50 85.50 90.50 75.50 79.00 62.00 91.50
s 25.00 66.50 50.00 82.50 43.00 66.00 10.50 69.00 35.50 74.50 63.50 91.50 57.50 78.50 13.50 85.00
- Druglikeness l 59.00 67.00 87.00 90.50 54.50 56.50 67.00 84.50 77.00 77.00 90.50 92.00 79.50 80.00 82.50 89.50
s 41.00 59.00 74.00 84.50 52.50 58.50 51.00 73.50 70.00 74.00 87.00 95.00 79.50 82.50 56.50 85.50
Average l 54.13 65.52 80.35 85.82 61.55 64.13 53.21 76.61 73.19 76.17 90.15 92.26 75.99 76.63 75.98 87.61
s 39.80 61.40 72.91 82.85 53.87 60.12 35.54 72.28 63.42 70.00 84.18 90.25 72.54 76.11 55.54 82.66
+ LogP & + TPSA l & l 34.00 56.50 54.50 68.00 50.50 54.50 27.50 54.00 47.00 63.00 59.50 84.00 47.50 63.00 30.50 79.50
s & s 25.50 51.00 43.50 62.50 33.50 37.50 22.50 53.00 40.00 64.00 56.50 83.00 50.00 68.00 14.00 78.50
+ LogP & - TPSA l & l 21.00 57.50 27.50 72.50 25.50 55.50 20.50 59.00 54.50 75.50 56.50 89.50 64.00 81.50 56.50 85.50
s & s 14.00 57.00 19.00 76.00 11.50 51.00 11.50 53.50 30.00 73.00 29.00 91.50 30.50 80.50 30.50 86.00
- LogP & + TPSA l & l 35.50 68.50 58.00 79.00 47.50 54.00 24.00 66.00 83.50 83.50 91.00 93.50 82.50 83.00 86.00 88.50
s & s 23.50 60.50 50.00 72.00 36.50 48.50 14.00 61.00 73.50 77.00 82.50 93.50 70.00 73.00 72.50 86.00
- LogP & - TPSA l & l 18.00 74.50 21.00 80.00 17.00 60.00 11.00 64.50 09.50 74.00 13.50 85.00 13.00 63.50 07.50 80.50
s & s 08.50 64.00 14.50 80.00 13.00 56.50 08.50 65.00 06.00 55.50 08.00 84.50 05.50 67.00 08.00 70.50
+ LogP & + HBDs l & l 26.50 52.50 52.00 64.50 41.50 46.00 21.00 55.50 49.50 67.00 60.50 84.00 45.50 66.00 14.50 76.00
s & s 22.50 52.00 43.00 65.50 37.50 50.00 11.50 56.50 38.50 71.00 47.00 81.50 34.00 59.00 05.50 77.50
+ LogP & - HBDs l & l 22.45 57.14 36.73 77.55 34.78 62.32 20.29 66.67 36.73 48.98 51.02 81.63 42.03 65.22 40.58 65.22
s & s 06.12 63.27 14.29 81.63 10.14 56.52 10.14 56.52 06.12 63.27 18.37 95.92 13.04 73.91 07.25 73.91
- LogP & + HBDs l & l 30.00 69.00 54.50 75.50 43.50 53.00 20.00 66.00 74.50 78.00 89.50 93.00 72.00 73.00 85.50 89.50
s & s 16.50 65.50 43.50 76.50 30.50 48.00 04.50 63.00 71.00 75.50 72.00 85.00 62.50 73.50 60.50 81.50
- LogP & - HBDs l & l 04.08 59.18 12.24 81.63 07.25 65.22 07.25 62.32 26.53 65.31 24.49 79.59 13.04 71.01 10.14 68.12
s & s 08.16 71.43 10.20 71.43 15.94 55.07 20.29 79.71 02.04 51.02 04.08 87.76 02.90 76.81 01.45 82.61
+ LogP & + HBAs l & l 29.00 53.50 45.50 65.00 35.50 43.50 20.00 53.00 50.00 65.50 52.00 76.00 44.00 60.50 34.00 73.00
s & s 24.50 56.50 41.50 58.50 26.50 39.50 15.50 62.00 37.00 56.00 56.00 85.00 44.00 62.50 12.50 73.50
+ LogP & - HBAs l & l 34.69 71.43 26.53 81.63 21.74 52.17 18.84 69.57 51.02 65.31 46.94 83.67 34.78 63.77 46.38 76.81
s & s 08.16 59.18 10.20 81.63 15.94 68.12 04.35 62.32 30.61 83.67 18.37 83.67 17.39 72.46 10.14 72.46
- LogP & + HBAs l & l 26.00 65.00 52.00 75.50 34.50 45.00 18.50 62.50 80.50 84.00 78.00 90.50 74.00 81.00 77.00 89.00
s & s 12.50 62.00 41.50 70.00 24.50 46.00 08.00 59.50 61.00 70.50 69.50 89.50 57.50 71.50 54.00 78.50
- LogP & - HBAs l & l 22.45 67.35 18.37 79.59 10.14 47.83 10.14 65.22 10.20 57.14 08.16 81.63 11.59 69.57 13.04 69.57
s & s 14.29 63.27 12.24 87.76 18.84 56.52 17.39 69.57 10.20 65.31 06.12 79.59 05.80 73.91 04.35 68.12
+ LogP & + Druglikeness l & l 27.50 58.00 42.50 72.00 37.50 47.50 18.00 60.50 40.00 78.50 46.50 86.00 49.50 75.00 27.00 84.00
s & s 08.00 68.00 31.00 73.00 19.50 47.00 02.00 58.00 17.50 82.50 23.50 85.00 26.00 74.00 04.00 84.00
+ LogP & - Druglikeness l & l 44.00 68.50 49.50 76.50 31.50 57.00 47.00 69.50 72.00 81.00 77.50 88.50 63.00 77.00 75.50 90.00
s & s 23.00 69.50 20.00 83.50 14.00 62.00 22.00 62.50 17.50 72.50 25.00 88.00 17.50 78.00 24.00 85.50
- LogP & + Druglikeness l & l 36.00 73.50 51.00 76.50 39.50 55.00 17.50 66.50 46.50 75.00 62.00 85.00 46.50 65.00 31.00 81.50
s & s 15.50 66.00 36.50 77.00 25.50 52.00 01.00 66.00 21.00 79.50 33.50 84.00 28.00 65.00 05.00 79.50
- LogP & - Druglikeness l & l 21.50 74.00 19.50 76.00 16.50 54.00 23.00 66.00 50.00 68.00 61.50 90.00 50.50 68.50 69.00 85.50
s & s 38.00 77.00 34.50 82.50 32.00 63.00 38.00 74.00 12.50 70.00 09.00 87.00 07.00 70.50 17.50 74.50
Average l & l 27.04 64.13 38.84 75.09 30.90 53.28 20.28 62.92 48.87 70.61 54.91 85.72 47.09 70.41 44.01 80.14
s & s 16.80 62.88 29.09 74.97 22.83 52.33 13.20 62.63 29.65 69.39 34.90 86.53 29.48 71.22 20.70 78.29
2.2 Refinement with $\text{Re}^{2}$
@@
After initialization, $\widehat{m}$ is immediately verified for its validity using an external toolkit, RDKit Landrum et al. (2013), a widely-used cheminformatics toolkit. If $\widehat{m}$ is chemically invalid, the concrete error message from RDKit is extracted as reliable domain feedback to inform the LLM $\mathcal{M}$ to correct $\widehat{m}$. Otherwise, $\text{Re}^{2}$ verifies whether $\widehat{m}$ meets the objective by checking whether $E(C_{p}(\widehat{m}))=1$, if not, its desired properties, $C_{p}(\widehat{m})$, are computed and compared to the original one, establishing reliable domain feedback, denoted fb, with correct direction and distance towards the objective, $|C_{p}(\widehat{m})-C_{p}(m)|$, to explicitly guide the LLM $\mathcal{M}$ to refine $\widehat{m}$. Additionally, as the previous ReDF, $\text{Re}^{2}$ retrieves from a prior database $\mathcal{D}$ a molecule $m_{e}$ that is similar to $\widehat{m}$ and meets the objective, serving as an example to enrich the domain feedback fb:
@@
where $\left\langle m_{e}^{\prime},\widehat{m}\right\rangle$ signifies the Morgan fingerprint-based Morgan (1965) Tanimoto similarity Bajusz et al. (2015) between a molecule $m_{e}^{\prime}$ in $\mathcal{D}$ and $\widehat{m}$. Subsequently, the described procedure is iterated for a pre-defined maximum number of iterations N or either can be stopped early if $\widehat{m}$ meets the objective, precisely depicted in Algorithm 1.